OpenStudy (anonymous):

Can you show me by steps how to solve this problem . ? y/3 + 8y/4 - 3/6= y + 7/2

OpenStudy (lalaly):

bring all the vaiables on one side \[\frac{y}{3}+\frac{8y}{4}-y = \frac{7}{2}+\frac{3}{6}\]

OpenStudy (lalaly):

\[\frac{y}{3}+2y-y=\frac{7}{2}+\frac{3}{6}\]

OpenStudy (lalaly):

\[\frac{y}{3}+y=\frac{7}{2}+\frac{3}{6}\]

OpenStudy (lalaly):

now u need a common denominator\[\frac{y}{3} +(y \times \frac{3}{3}) = \frac{7}{2} +\frac{1}{2}\]

OpenStudy (lalaly):

\[\frac{y}{3} + \frac{3y}{3} = \frac{8}{2}\]

OpenStudy (lalaly):

\[\frac{4y}{3}=4\]

OpenStudy (lalaly):

\[\cancel{3} \times \frac{4y}{\cancel{3}} = 4 \times 3\]

OpenStudy (lalaly):

\[4y=12\]\[y=\frac{12}{4} = 3\]

OpenStudy (anonymous):

thank youu !!

OpenStudy (lalaly):

your welcome

OpenStudy (anonymous):

ii may need more help tho . ! :D

OpenStudy (lalaly):

hehe :D ill be glad to help

OpenStudy (anonymous):

well , my daughter is taking a 25 question test ; maybe we can review her work ?

OpenStudy (anonymous):

write and solve the algebraic inequality ; -5 and a number is at least 40.

OpenStudy (anonymous):

-40 *

OpenStudy (anonymous):

this problem has answer choices , && -35 isn't on there . It's a little hard too explain , shall I just ask another question . ?

OpenStudy (anonymous):

 Here is the next problem.